# dot-PE: Sampler-free gravitational wave inference using matrix multiplication

### v1

@@ Abstract
Parameter estimation (PE) for compact binary coalescence (CBC) events observed by gravitational wave (GW) laser interferometers is a core task in GW astrophysics. We present a method to compute the posterior distribution efficiently without relying on stochastic samplers. First, we show how to select sets of intrinsic and extrinsic parameters that efficiently cover the relevant phase space. We then show how to compute the likelihood for all combinations of these parameters using dot products. We describe how to assess and tune the integration accuracy, making the outcome predictable and adaptable to different applications. The low computational cost allows full PE in minutes on a single CPU, with the potential for further acceleration using multiple CPUs or GPUs. We implement this method in the dot-PE [^1] https://github.com/JonathanMushkin/dot-PE package, enabling sensitive searches using the full evidence integral for precessing CBCs and supporting large waveform banks ($\sim 10^{6}$ waveforms), regardless of waveform generation cost.
@@ 1 Introduction
Within the Bayesian parameter estimation framework, the process of inferring the physical properties of the emitting systems from the GW events is laborious. The main challenge is to efficiently explore the 15-dimensional parameter space of the waveform. Numerous sampling techniques and codes have been developed to address it (including, but not limited to, [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]). For a recent review, see [27].
@@ 2 Summary of this work
Figure 1: Flowchart of dot-PE. Intrinsic waveform bank, (block 1, Section 4.3) is performed once per mass range and can be re-used for different events. Block 2 is the output of a conventional search pipeline: detector strains and trigger information. In block 3, the phase space is reduced. The trigger is assigned to a bank by its ${\mathcal{M}}$. Intrinsic samples are accepted based on their incoherent maximum-likelihood (Section 4.4, reducing the samples from $N_{\rm int.}$ to $N_{\rm int.}^{\prime}$ (see Section 5.1 and Fig. 5 for the relation between $N_{\rm int.}^{\prime}$ and the expected performance of the integration). Extrinsic samples are drawn using adaptive-importance sampling conditioned on the data and a subset of $N_{c}$ intrinsic samples (Section 4.5). In block 4, the intrinsic and extrinsic components are combined to evaluate the inner products and likelihoods, per intrinsic-extrinsic-phase combination (Section 4.2). The outputs of the method are the evidence integral (block 5) and samples drawn from the posterior distribution (block 6).
@@
Naively, the priors set parameter ranges too wide to cover properly with reasonable amount of samples. We therefore use the data to narrow the ranges:
@@
    The pipeline trigger itself carries a rough estimate for the chirp-mass, limiting the intrinsic-parameter samples to one particular pre-computed bank of waveforms.
@@
    Given the data, we include or exclude each of the $N_{\rm int.}$ intrinsic waveforms from the bank according to their likelihood, incoherently maximized across detectors, eventually keeping $N_{\rm int.}^{\prime}$.
@@
The luminosity-distance $d_{L}$ will be marginalized over. Once the components for each point are calculated ($A_{ix}=A_{f}\big{(}\bm{\theta}_{\rm int}^{(i)}\big{)}$, $B_{ef}=B_{f}\big{(}\bm{\theta}_{\rm ext}^{(e)}\big{)}$,and $C_{of}=C_{f}\big{(}\phi_{\rm ref}^{(o)}\big{)}$), the distance-marginalized likelihood $\overline{\mathcal{L}}$ at all possible combinations can be evaluated through a series of matrix multiplications, and a distance marginalization function $F$:
@@
with $w$ being the probabilistic weights of the samples, related to the drawing process. The Bayesian evidence is the optimal test statistic for a search [28]. The applications of this method are numerous:
@@
    Detection: The main motivation is to compute $\mathcal{Z}$ quickly and reliably enough to the point it can be used as a test statistic for a search, as a second-stage after the more traditional match-filtering pipelines.
@@
    Posterior samples: The summand of Eq. (3) are the probabilistic weights of individual points, which allows for drawing samples for parameter estimation. By modifying the weights $w$ in post-processing, we can resample from different priors.
@@
    The individual maximum-likelihood point can be used for fitting, for signal consistency checks.
@@
Figure 2: Convergence of the evidence integral across intrinsic waveform banks with varying mass ranges (Section 4.3), shown in different colors and markers as indicated in the top right legend. Left: Absolute value of the relative error $|\ln\mathcal{Z}-\ln\hat{\mathcal{Z}}|$ for $10242^{10}=$ 1024 injections per mass range, with $N_{\rm int.}$ from $2^{6}$ to $2^{16}\approx 6.5\times 10^{4}$, and fixed $N_{\rm ext.}=2^{10}$. Reference values $\hat{\mathcal{Z}}$ are computed with much finer banks of $N_{\rm int.}=2^{18}\approx 2.6\times 10^{5}$. Solid lines show medians; shaded bands span 25th–75th percentiles. The error decreases with $N_{\rm int.}$ until saturation due to extrinsic sample count and $\phi_{\rm ref}$ resolution. Center: Same as left, but plotted against $N_{\rm int.}^{\prime}$, the number of intrinsic samples retained after pre-selection (Section 4.4). The improved alignment between mass ranges—relative to the left panel—illustrates that convergence is largely governed by the number of relevant samples near the posterior support. Right: Absolute value of the relative error versus $N_{\rm ext.}$, ranging from $2$ to $2^{10}$, for fixed $N_{\rm int.}=2^{16}$.
@@ 3.1 Bayesian Inference
[…]
@@ 3.2 Integration methods
[…]
@@
[…]
@@
respectively. The probabilistic weights enable us to resample, that is, to draw samples from the $N$ according to the probabilistic weights $\left\{p_{n}\right\}_{n=1}^{N}$. The samples can be used to describe the posterior, visualize it using one and two dimensional histograms, and compute arbitrary integrals with the posterior as a measure. The distribution of the weights reveals the confidence in the integral estimate, through the effective sample size [31, 32]:
@@
which satisfies $1\leq N_{\rm eff.}\leq N$. After computing $\{p_{n}\}_{n=1}^{N}$, the effective sample size can indicate whether the integral is dominated by a single or a few samples, thus prone to large variation, or comprised of many similar contributions, hence more reliable  [^3] While the effective sample size is justified for i.i.d samples, we expect the same qualitative behavior in QMC samples..
@@ 4.1 Waveform Decomposition
Combining all the effects, the strain at detector $h$ due to the gravitational waves is
@@ 4.2 Likelihood components as matrix products
$\ln\overline{\mathcal{L}}_{ieo}=\int{\rm d}d_{L}\Pi(d_{L})\exp\left(\frac{\langle d\mid h\rangle_{ieo}}{d_{L}/d_{L}^{0}}-\frac{1}{2}\frac{\langle h\mid h\rangle_{ieo}}{\left(d_{L}/d_{L}^{0}\right)^{2}}\right)$ (38)
The distance marginalization in this work is done using the cogwheel distance marginalization routines [25]. It does not benefit from the matrix form of the Eqs. (34) and (35). It is therefore beneficial to remove $(i,e,o)$ combinations have small contribution to $\mathcal{Z}$ or little agreement with the data. Thus, we evaluate Eq. (38) only for samples with
@@ 4.3 Intrinsic waveform banks
The sampling distribution $\pi^{\prime}(\bm{\theta}_{\rm int})$ used to draw the waveform bank differs from the assumed physical prior. Lower masses perform more cycles in band, so sampling is done uniformly in $\ln{\mathcal{M}}$. Lower mass ratios exhibit stronger higher-order modes, so sampling is done uniformly in $\ln q$. $\theta_{JN}$ is sampled uniformly between 00 and $\pi$, and the spin priors remain unchanged. The probability of each intrinsic parameters sample is reweighted by
@@ 4.4 Intrinsic Waveforms Pre-selection
$\ln\mathcal{L}^{\rm inco.\,ML}_{i}$ $=\sum_{k}\max_{i,o}\ln\mathcal{L}^{\rm ML}_{ikot}$ (45)
$\geq\max_{i}\ln\mathcal{L}_{i}^{\rm inco.ML}-\Delta\mathcal{L}^{\rm inco.\,ML}$
with $\Delta\mathcal{L}^{\rm inco.\,ML}=20$. In its current implementation, the runtime of dot-PE is dominated by the pre-selection stage, as can be seen in Appendix D.
@@ 4.5 Extrinsic samples
Given data from an event, the phase space of possible extrinsic parameters is much larger than what the data allows, due to the small correlation timescale of the signals, and the geometric relation between sky-positions and signal arrival time at the individual detectors (Eq. (24)). Therefore we adopt the adaptive-importance sampling method presented in [25] (described briefly in Appendix C) to draw extrinsic samples in an informed way. Given intrinsic parameters and the data, this method produces samples of $t_{\oplus}$ and $\bm{\hat{n}}$ based on estimates of the times of arrival of the signal at each detector. $\psi$ is drawn uniformly from $(0,\pi)$. The method also allows us to evaluate the marginalized likelihood ¯$\overline{\mathcal{L}}(\bm{\theta}_{\rm int})$ per choice of intrinsic parameters, through a weighted ensemble average over $N_{\rm ext.marg}$ extrinsic samples, trapezoidal quadrature over $\phi_{\rm ref}$, and analytic integration over $d_{L}$. For our purposes, we want to draw a single set of extrinsic samples, decoupled from intrinsic samples. This is done by aggregating extrinsic samples drawn using $N_{c}=16$ different intrinsic samples. This aggregation is motivated by previous works that have shown correlations between the inclination angle (treated here as intrinsic) and sky location [53, 52]. After constructing the aggregated ensembles, the proposal distributions $\pi^{\prime}$ are combined (as in Eq. (30) of [25]), and importance-sampling weights are assigned to each extrinsic sample as
@@
where ¯$\overline{\mathcal{L}}_{ie}$ is the distance and phase marginalized likelihood, for a given intrinsic-extrinsic sample pair. Under one simplistic limiting case, the likelihood per extrinsic sample is the same for all intrinsic samples used in the integral ¯$\overline{\mathcal{L}}_{ie}=\overline{\mathcal{L}}_{e}$. The effective sample size can be factorized
@@
and the right-most ratio is the same in form as the effective sample size evaluated per proposal in [25]. We therefore require that each proposal (conditioned on $\bm{\theta}_{\rm int}$) satisfies $N_{\rm eff.}\geq 100$. In the opposite limiting case, the rows of ¯$\overline{\mathcal{L}}_{ie}$ are uncorrelated, and we can replace ¯$\overline{\mathcal{L}}_{ie}$ and ¯$\overline{\mathcal{L}}_{ie}^{2}$ with their expectation values (using $\mathbb{E}$ to differentiate from the distance and phase marginalized value)
@@ 5 Performance
We test the performance of our method against injections into synthetic noise realizations, using the Hanford–Livingston–Virgo detector network with O3 sensitivities. 1,024 injections are performed for each of 6 chirp-mass ranges: $(3,3.2){\rm M}_{\odot}$, $(5,5.6){\rm M}_{\odot}$, $(10,12){\rm M}_{\odot}$, $(20,30){\rm M}_{\odot}$, $(50,100){\rm M}_{\odot}$, and $(100,300){\rm M}_{\odot}$. The injection prior are the same as our assumed physical priors, as used in the intrinsic bank creation and in extrinsic sample drawing: uniform ${\mathcal{M}}$ in range, uniform $\ln q$ restricted to $q\in(0.2,1)$, uniform ${\chi_{\rm eff}}\in(-1,1)$, uniform $C^{\perp}$ and isotropic in-plane spin for each binary component, isotropic sky position, uniform polarization angle and uniform luminosity–volume prior ($\pi(d_{L})\propto d_{L}^{2}$). See [52] for definitions of the parameters. We impose on the injections $\langle h\mid h\rangle\in(70,200)$ appropriate for most detections. We also perform a small number of individual injections, to validate the performance of dot-PE for parameter estimation against nested sampling using cogwheel, see Section 5.2.
@@ 5.1 Integration
We set a target accuracy of ±$\pm 1$ in evaluating $\ln\mathcal{Z}$ using our work. In the Gaussian-posterior limit, $\ln\mathcal{Z}\propto\rho^{2}/2+\mathcal{O}(\ln\rho)$. Although the detection limit could depend on the mass, it is roughly set to $\rho=8$. The distance and $\rho^{2}$ are related through (following Eqs. (36)–(37)):

### v3

# Sampler-free gravitational wave inference using matrix multiplication

@@ Abstract
Parameter estimation (PE) for compact binary coalescence (CBC) events observed by gravitational wave (GW) laser interferometers is a core task in GW astrophysics. We present a method to compute the posterior distribution efficiently without relying on stochastic samplers. First, we show how to select sets of intrinsic and extrinsic parameters that efficiently cover the relevant phase space. We then show how to compute the likelihood for all combinations of these parameters using dot products. We describe how to assess and tune the integration accuracy, making the outcome predictable and adaptable to different applications. The low computational cost allows full PE in minutes on a single CPU, with the potential for further acceleration using multiple CPUs or GPUs. We implement this method in the dot-PE [^1] https://github.com/JonathanMushkin/dot-PE package, enabling sensitive searches using the full evidence integral for precessing CBCs and supporting large waveform banks ($\sim 10^{5}$$10^{6}$ waveforms), regardless of waveform generation cost.
@@ 1 Introduction
Within the Bayesian parameter estimation framework, the process of inferring the physical properties of the emitting systems from the GW events is computationally expensive. The main challenge is to efficiently explore the 15-dimensional parameter space of the waveform. Numerous sampling techniques and codes have been developed to address it (including, but not limited to, [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]). For a recent review, see [27].
@@ 2 Summary of this work
Figure 1: Flowchart of dot-PE. Intrinsic waveform bank, (block 1, Section 4.3) is performed once per mass range and can be re-used for different events. Block 2 is the output of a conventional search pipeline: detector strains and trigger information. In block 3, the phase space is reduced. The trigger is assigned to a bank by its $\mathcal{M}$}. Intrinsic samples are accepted based on their incoherent maximum-likelihood (Section 4.4), reducing the samples from $N_{\rm int.}$ to $N_{\rm int.}^{\prime}$ (see Section 5.1 and Fig. 5 for the relation between $N_{\rm int.}^{\prime}$ and the expected performance of the integration). Extrinsic samples are drawn using adaptive-importance sampling conditioned on the data and a subset of $N_{c}$ intrinsic samples (Section 4.5). In block 4, the intrinsic and extrinsic components are combined to evaluate the inner products and likelihoods, per intrinsic-extrinsic-phase combination (Section 4.2). The outputs of the method are the evidence integral (block 5) and samples drawn from the posterior distribution (block 6).
@@
Traditional matched-filtering pipelines (e.g. cite [28, 29, 30, 31]) search for candidate signals by performing matched filtering (cross-correlation) of the detector strain against a bank of template waveforms. Times at which a particular template yields a large signal-to-noise (SNR) value are reported as triggers. Given such a trigger, the goal of this work is to evaluate the likelihood with sufficient density across the entire parameter space. Naively, the prior ranges are too wide to be covered with a reasonable number of samples. We therefore narrow the parameter space using the following steps:
@@
[…]
@@
    Given the data, we include or exclude each of the $N_{\rm int.}$ intrinsic waveforms according to a likelihood maximized incoherently across detectors (i.e., the maximization over $\theta_{\rm ext}$, $\phi_{\rm ref}$, and $d_{L}$ is done per detector, without cross-detector consistency). This pre-selection retains $N_{\rm int.}^{\prime}$ intrinsic samples (Sec. 4.4).
@@
The luminosity-distance $d_{L}$ will be marginalized over. Once the components for each point are calculated ($A_{if}=A_{f}\big{(}\bm{\theta}_{\rm int}^{(i)}\big{)}$, $B_{ef}=B_{f}\big{(}\bm{\theta}_{\rm ext}^{(e)}\big{)}$,and $C_{of}=C_{f}\big{(}\phi_{\rm ref}^{(o)}\big{)}$), the distance-marginalized likelihood $\overline{\mathcal{L}}$ at all possible combinations can be evaluated through a series of matrix multiplications, and a distance marginalization function $F$:
@@
with $w$ being the probabilistic weights of the samples, related to the drawing process. The applications of this method are numerous:
@@
    Detection: The main motivation is to compute $\mathcal{Z}$, which is the optimal test statistic for a search [32], quickly and reliably, as a second stage after a matched-filtering pipeline.
@@
    Parameter estimation: Detected events are analyzed by drawing samples from the posterior distribution of the source parameters, defined by the probabilistic weights in Eq. (3). By modifying the weights $w$ in post-processing, we can also resample from different priors.
@@
    The individual maximum-likelihood point can be used for fitting and signal consistency checks [33].
@@
[…]
@@ 3.1 Bayesian Inference
The inner product $\langle\cdot\mid\cdot\rangle$ is the sum of detector-wise frequency integral, approximated by a sum:
@@ 3.2 Integration methods
$\mathcal{Z}\approx\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\bm{\theta}_{i}),\quad\bm{\theta}_{i}\sim\pi(\bm{\theta})\,\text{ i.i.d}.$ (14)
@@
$\mathcal{Z}\approx\frac{1}{N}\sum_{i=1}^{N}\frac{\pi(\bm{\theta}_{i})}{\pi^{\prime}(\bm{\theta}_{i})}\mathcal{L}(\bm{\theta}_{i}),\quad\bm{\theta}_{i}\sim\pi^{\prime}(\bm{\theta})\,\text{ i.i.d}.$ (15)
@@
respectively. The probabilistic weights enable us to resample, that is, to draw from the $N$ samples according to the probabilistic weights $\left\{p_{n}\right\}_{n=1}^{N}$. The samples can be used to describe the posterior, visualize it using one- and two-dimensional histograms, and compute integrals under the posterior distribution, such as mean values, variances, and the probabilities of lying above or below particular values. The distribution of the weights quantifies the confidence in the evidence integral estimate, through the effective sample size [36, 37]:
@@
which satisfies $1\leq N_{\rm eff.}\leq N$. After computing $\{p_{n}\}_{n=1}^{N}$, the effective sample size can indicate whether the evidence estimate is dominated by a single or a few samples, making it prone to large variation, or comprises many similar contributions, making it more reliable  [^3] While the effective sample size is justified for i.i.d samples, we expect the same qualitative behavior in QMC samples..
@@
A dominant failure mode of both $MC$ and NS is multimodality, i.e. the presence of disjoint posterior peaks in the parameter space. Even though modern implementations attempt to mitigate this issue, it remains a limiting factor for both methods [43, 41, 44].
@@ 4.1 Waveform Decomposition
Combining all the effects, the strain due to the gravitational waves, as measured at the detector, is
@@ 4.2 Likelihood components as matrix products
$\ln\overline{\mathcal{L}}_{ieo}=\int{\rm d}d_{L}\pi(d_{L})\exp\left(\frac{\langle d\mid h\rangle_{ieo}}{d_{L}/d_{L}^{0}}-\frac{1}{2}\frac{\langle h\mid h\rangle_{ieo}}{\left(d_{L}/d_{L}^{0}\right)^{2}}\right),$ (38)
where $\pi(d_{L})$ distance prior. The distance marginalization in this work is done using the cogwheel distance marginalization routines [25]. It does not benefit from the matrix form of the Eqs. (34) and (35). It is therefore beneficial to remove $(i,e,o)$ combinations that have small contribution to $\mathcal{Z}$ or little agreement with the data. Thus, we evaluate Eq. (38) only for samples with
@@ 4.3 Intrinsic waveform banks
The sampling distribution $\pi^{\prime}(\bm{\theta}_{\rm int})$ used to draw the waveform bank differs from the assumed physical prior. Lower masses perform more cycles in band, so sampling is done uniformly in $\ln{\mathcal{M}}$. Lower mass ratios exhibit stronger higher-order modes, so sampling is done uniformly in $\ln q$. $\theta_{JN}$ is sampled uniformly between 0 and $\pi$, and the spin priors remain unchanged. The probability of each intrinsic parameters sample is reweighted by
@@ 4.4 Intrinsic Waveforms Pre-selection
The selection criterion on the incoherently maximized likelihood is
$\ln\mathcal{L}^{\rm inco.\,ML}_{i}$ $=\sum_{k}\max_{o,t}\ln\mathcal{L}^{\rm ML}_{ikot}$ (45)
$\geq\max_{j}\ln\mathcal{L}^{\rm inco.\,ML}_{j}-\Delta\mathcal{L}^{\rm inco.\,ML},$
with $\Delta\mathcal{L}^{\rm inco.\,ML}=20$ in this work. The number of intrinsic samples retained after the procedure is denoted $N_{\rm int.}^{\prime}$. In its current implementation, the runtime of dot-PE is dominated by this pre-selection stage (see Appendix D).
@@ 4.5 Extrinsic samples
Given data from an event, the phase space of possible extrinsic parameters is much larger than what the data allows, due to the small correlation timescale of the signals, and the geometric relation between sky-positions and signal arrival time at the individual detectors (Eq. (24)). Therefore we adopt the adaptive-importance sampling method presented in [25] (described briefly in Appendix C) to draw extrinsic samples in an informed way. Given intrinsic parameters and the data, this method produces samples of $t_{\oplus}$ and $\bm{\hat{n}}$ based on estimates of the times of arrival of the signal at each detector. $\psi$ is drawn uniformly from $(0,\pi)$. The method also allows us to evaluate the marginalized likelihood $\overline{\mathcal{L}}(\bm{\theta}_{\rm int})$ per choice of intrinsic parameters, through a weighted ensemble average over $N_{\rm ext.marg}$ extrinsic samples, trapezoidal quadrature over $\phi_{\rm ref}$, and analytic integration over $d_{L}$. For our purposes, we want to draw a single set of extrinsic samples, decoupled from intrinsic samples. This is done by aggregating extrinsic samples drawn using $N_{c}=16$ different intrinsic samples. This aggregation is motivated by previous works that have shown correlations between the inclination angle (treated here as intrinsic) and sky location [60, 59]. After constructing the aggregated ensembles, the proposal distributions $\pi^{\prime}$ are combined (as in Eq. (30) of [25]), and importance-sampling weights are assigned to each extrinsic sample as
@@
where $\overline{\mathcal{L}}_{ie}$ is the distance and phase marginalized likelihood, for a given intrinsic-extrinsic sample pair. Under one simplistic limiting case, the likelihood per extrinsic sample is the same for all intrinsic samples used in the integral $\overline{\mathcal{L}}_{ie}=\overline{\mathcal{L}}_{e}$. The effective sample size can be factorized
@@
and the right-most ratio is the same in form as the effective sample size evaluated per proposal in [25]. We therefore require that each proposal (conditioned on $\bm{\theta}_{\rm int}$) satisfies $N_{\rm eff.}\geq 100$. In the opposite limiting case, the rows of $\overline{\mathcal{L}}_{ie}$ are uncorrelated, and we can replace $\overline{\mathcal{L}}_{ie}$ and $\overline{\mathcal{L}}_{ie}^{2}$ with their expectation values (using $\mathbb{E}$ to differentiate from the distance and phase marginalized value)
@@ 5 Performance
We test the performance of our method against injections into synthetic noise realizations, using the Hanford–Livingston–Virgo detector network with O3 sensitivities. 1,024 injections are performed for each of 6 chirp mass ranges: $(3,3.2){\rm M}_{\odot}$, $(5,5.6){\rm M}_{\odot}$, $(10,12){\rm M}_{\odot}$, $(20,30){\rm M}_{\odot}$, $(50,100){\rm M}_{\odot}$, and $(100,300){\rm M}_{\odot}$. The injection prior are the same as our assumed physical priors, as used in the intrinsic bank creation and in extrinsic sample drawing: uniform $\mathcal{M}$} in range, uniform $\ln q$ restricted to $q\in(0.2,1)$, uniform ${\chi_{\rm eff}}\in(-1,1)$, uniform $C^{\perp}$ and isotropic in-plane spin for each binary component, isotropic sky position, uniform polarization angle and uniform luminosity–volume prior ($\pi(d_{L})\propto d_{L}^{2}$). See [59] for definitions of the parameters. We impose on the injections $\langle h\mid h\rangle\in(70,200)$ appropriate for most detections. We also perform a small number of individual injections, to validate the performance of dot-PE for parameter estimation against nested sampling using cogwheel, see Section 5.2.
@@ 5.1 Integration
We set a target accuracy of $\pm 1$ in evaluating $\ln\mathcal{Z}$ using our work. In the Gaussian-posterior limit, $\ln\mathcal{Z}\propto\rho^{2}/2+\mathcal{O}(\ln\rho)$. Although the detection limit could depend on the mass, it is roughly set to $\rho=8$. The distance and $\rho^{2}$ are related through (following Eqs. (36)–(37)):